# Computation Mechanism Behind LLM Position Generalization

### v1

@@ 2.1 Self-Attention and Positional Encoding
${\bm{o}}_{i}=\sum_{j\leq i}\frac{w({\bm{q}}_{i},{\bm{k}}_{j},i-j)}{\sum_{j^{\prime}\leq i}\exp{w({\bm{q}}_{i},{\bm{k}}_{j^{\prime}},i-j^{\prime})}}\bm{v}_{j}$ (1)
. In spite of the existence of other choices of function $w(\cdots)$ like Alibi Press et al. (2021), the de-facto mainstream choice is RoPE Su et al. (2021). It decomposes ${\bm{q}}$ and ${\bm{k}}$ vectors into 2-D tuples and lets them rotate in angle $(i-j)\theta_{r}$, where each 2-D tuple $r$ has a different rotating “angular speed” $\theta_{r}$.
@@ 3 LLMs Disentangle Position and Semantics in Attention
Counter-intuitively, in this section, we reveal that LLMs learn a special feature pattern to empirically simplify the logit function $w(\cdots)$. The resulting attention logits can be approximately disentangled as an arithmetic addition of position relevance (determined by $i-j$ and ${\bm{q}}_{i}$) and semantic importance (determined by ${\bm{k}}_{j}$). We will start with an interesting observation of low-rank components in attention maps in Sec 3.1. Taking it as an inspiration, Sec 3.2 shows how the attention logits can be approximately disentangled into position and semantic-related components. Finally, Sec 3.3 shows how LLMs computationally achieve this mechanism by enforcing a special pattern in key and query vectors.
@@ 3.2 The Disentanglement Law of Attention Logits
Therefore, instead of studying the real attention logits, we use a fake distance value $d$ to replace the real distance $i-j$: $w(d,{\bm{q}}_{i},{\bm{k}}_{j})$. In light of the previous discussion, we fix a query vector ${\bm{q}}$ and visualize the following fake logit matrix $W^{\prime}\in{\mathbb{R}}^{n\times n}$ where $W_{d,j}^{\prime}=w(d,{\bm{q}},{\bm{k}}_{j})$ in Fig 3(a). After this substitution, the new matrix $W^{\prime}$ shows apparent vertical and horizontal patterns, suggesting prominent distance-wise and key-wise components.
@@
. Essentially, ${\bm{a}}$ and ${\bm{b}}$ are the average column and row of $W^{\prime}$, respectively, with a constant offset of $-\frac{1}{2n^{2}}\sum_{d^{\prime},j^{\prime}}W_{d^{\prime},j}^{\prime}$. This disentanglement process is visualized in Fig 3(b), where the key-axis and distance-axis components at two corners align with the patterns of the original fake logit matrix well. Once we combine these two components, their summation again demonstrates high similarity with the original matrix with details. This approximation has a correlation coefficient of 0.9470, explaining the vast majority of the logits’ variance by the two simple 1-dimensional components. We list more examples of such approximation in Fig 3(c) and Appendix D. These results indicate an approximated disentanglement of attention logits between positional relevance and semantic relevance:
@@ 4.1 Tolerance to Position Perturbations
The results, presented in Fig. 5, show that LLMs exhibit robustness to these perturbation methods. The original model has a perplexity of 3.688 on the ArXiv documents Gao et al. (2020). Text transposition has a minimal impact on perplexity, with only a 0.02 increase when 1% of tokens are shuffled up to a distance of 1000 tokens. This suggests that LLMs do not rigidly depend on strict word order. Our intervention techniques simulate this linguistic transposition effect in Fig. 5(b) and  5(c). Feature transposition also introduces a modest increase in perplexity, indicating that while position indices contribute to contextual representation, their precise ordering is not always critical in each self-attention layer. As further analysis, when the position encoding contains perturbed indices, the perplexity still has a marginal increase when 10% of token positions were perturbed by ±$\pm 10$, or increase by an absolute value of 0.01 when 1% of tokens has position encoding perturbed by ±$\pm 10$. These phenomena further align with the previous observations that position information acts as a disentangled additive factor rather than being tightly entangled with semantic relationships.
@@ 4.2 When and How LLMs Generalize to Longer Texts
Recent techniques like LM-Infinite and InfLLM enable LLMs to generalize to longer text sequences than those encountered during training. The common practice adopted by these techniques is to modify the relative position before applying the original self-attention mechanism. This is equivalent to applying self-attention over a modified $[{\bm{k}}_{i}]$ and $[{\bm{v}}_{i}]$ sequence with probably, which might be different (sometimes significantly shorter) than the original $[{\bm{k}}_{i}]$ and $[{\bm{v}}_{i}]$ sequences. More specifically, in those techniques, the resulting sequence of features usually appears as if they are of the following positional distances:
@@
This is in contrast to the intuition we obtained from common machine learning practices: why do LLMs train purely from shorter texts that generalize to extreme lengths (e.g., 200M in LM-Infinite) with only minor modifications to the model architecture? Moreover, little explicit design was implemented in modern SotA LLMs to enable this extreme generalization. This phenomenon could find support in our analysis: even though posed to unseen extreme lengths, LLMs do not bind positional relevance information with the semantic features of the contextual tokens. In other words, the ${\bm{k}}$ and ${\bm{v}}$ vectors could be interpreted as a pool of semantic features. The self-attention mechanism approximately and additively applies the position component to the pool. As long as the resulting distance sequence is similar to its normal shape $[L_{PT},L_{PT}-1,\ldots,1,0]$, the attention output vector will reside in its normal distribution. Therefore, technically, the attention output vector is still in-distribution, so the remaining parts of LLMs on top of the attention outputs will not take out-of-distribution features as inputs.
To verify this claim, we visualize the attention output vectors of an arbitrary layer using the technique above. This is a projection down to a 2-D plane using PCA [^4] https://scikit-learn.org/stable/modules/generated/sklearn.decomposition.PCA.html. The blue dots are the normal attention output vectors, which mark their normal distribution. Then, we select a set of ${\bm{q}}$ vectors and associate them with different colors. For each vector, we apply it over a sub-sequence of length $L_{PT}$: $[{\bm{k}}_{i},{\bm{k}}_{i+1},\ldots,{\bm{k}}_{i+L_{PT}}]$ and $[{\bm{v}}_{i},{\bm{v}}_{i+1},\ldots,{\bm{v}}_{i+L_{PT}}]$. As we vary the value of starting position $i$, we trace the output vector with colored broken lines. As shown in Figure 6, these lines, though extending to different directions and different ranges depending on the ${\bm{q}}$, still wander within the range of normal attention output vector distribution. This further validates our explanation of the length generalization and provides insights for future manipulation of the self-attention module for research purposes.
@@ 5 Conclusions and Future Work
In this work, we investigated the computational mechanisms behind the position generalization capabilities of LLMs. We first demonstrated that attention logits in LLMs can be approximately disentangled into independent components representing positional and semantic relevance. This finding suggests a structured decomposition within the model’s internal computations. Through empirical analysis, we further examined various forms of position generalization at the LLM representation level. These insights provide both computational explanations and insights into controlling these phenomena. Future research could delve deeper into the specific architectural choices and training data patterns that contribute to this robustness.

### v2

@@ 2.1 Self-Attention and Positional Encoding
${\bm{o}}_{i}=\sum_{j\leq i}\frac{w({\bm{q}}_{i},{\bm{k}}_{j},i-j)}{\sum_{j^{\prime}\leq i}\exp{w({\bm{q}}_{i},{\bm{k}}_{j^{\prime}},i-j^{\prime})}}\bm{v}_{j}.$ (1)
In spite of the existence of other choices of function $w(\cdots)$ like Alibi Press et al. (2021), the de-facto mainstream choice is RoPE Su et al. (2021). It decomposes ${\bm{q}}$ and ${\bm{k}}$ vectors into 2-D tuples and lets them rotate in angle $(i-j)\theta_{r}$, where each 2-D tuple $r$ has a different rotating “angular speed” $\theta_{r}$.
@@ 3 LLMs Disentangle Position and Semantics in Attention
Counter-intuitively, in this section, we reveal that LLMs learn a special feature pattern to empirically simplify the logit function $w(\cdots)$. The resulting attention logits can be approximately disentangled as an arithmetic addition of position relevance (determined by $i-j$ and ${\bm{q}}_{i}$) and semantic importance (determined by ${\bm{k}}_{j}$). We will start with an interesting observation of low-rank components in attention maps in Sec 3.1. Taking it as an inspiration, Sec 3.2 shows how the attention logits can be approximately disentangled into position and semantic-related components. Sec 3.3 generalizes the discussion to the relative position encoding scheme in general. Finally, Sec 3.4 shows how LLMs computationally achieve this mechanism by enforcing a special pattern in key and query vectors.
@@ 3.2 The Disentanglement Law of Attention Logits
Therefore, instead of studying the real attention logits, we use a fake distance value $d$ to replace the real distance $i-j$: $w(d,{\bm{q}}_{i},{\bm{k}}_{j})$. In light of the previous discussion, we fix a query vector ${\bm{q}}$ (for which we use the last token in the document to maximize the attention scope) and visualize the following fake logit matrix $W^{\prime}\in{\mathbb{R}}^{n\times n}$ where $W_{d,j}^{\prime}=w(d,{\bm{q}}_{i},{\bm{k}}_{j})$ in Fig 3(a). After this substitution, the new matrix $W^{\prime}$ shows apparent vertical and horizontal patterns, suggesting prominent distance-wise and key-wise components.
@@
$\begin{split}a_{d}&=\frac{1}{n}\sum_{j^{\prime}}W_{d,j^{\prime}}^{\prime}-\frac{1}{2n^{2}}\sum_{d^{\prime},j^{\prime}}W_{d^{\prime},j^{\prime}}^{\prime}\\ b_{j}&=\frac{1}{n}\sum_{d^{\prime}}W_{d^{\prime},j}^{\prime}-\frac{1}{2n^{2}}\sum_{d^{\prime},j^{\prime}}W_{d^{\prime},j^{\prime}}^{\prime}.\\ \end{split}$ (3)
Essentially, ${\bm{a}}$ and ${\bm{b}}$ are the average column and row of $W^{\prime}$, respectively, with a constant offset of $-\frac{1}{2n^{2}}\sum_{d^{\prime},j^{\prime}}W_{d^{\prime},j}^{\prime}$. This disentanglement process is visualized in Fig 3(b), where the key-axis and distance-axis components at two corners align with the patterns of the original fake logit matrix well. Once we combine these two components, their summation again demonstrates high similarity with the original matrix with details. This approximation has an average correlation coefficient of 0.9470 across layers, with a minimum linear correlation of the disentangling approximation of 0.914 and a maximum of 0.967, explaining the vast majority of the logits’ variance by the two simple 1-dimensional components. We list more examples of such approximation in Fig 3(c) and Appendix D. These results indicate an approximated disentanglement of attention logits between positional relevance and semantic relevance:
@@
Discussion: Triviality of Eq. 4
We here argue that RoPE does not inherently disentangle positional and semantic information by design. As an empirical verification, we additionally experiment on a randomly initialized Llama-3.2-3B model. Using the same setup as described before, we carry out a 3-axis linear approximation of the logit map in this model on an input of length 4096. In this case, we only observe a 0.028 linear correlation coefficient between the original and approximated attention logits. RoPE is mathematically equivalent to evaluating an inverse discrete Fourier transform along position $d$, which can express arbitrarily complex correlations between ${\bm{q}},{\bm{k}},d$ as shown in the following Theorem, with a proof listed in Appendix E.
Theorem 1.
Let $f(d,x,y)\in[-C,C]$ be any bounded function over relative position $d\in\mathbb{Z}$ and semantic inputs $x\in\mathcal{X},y\in\mathcal{Y}$ with smoothness properties[6]. Then, for sufficiently large feature dimension $n=2N$, there exists a set of rotation frequencies $\{\theta_{r}\}$, and features $\bm{k}(x),\bm{q}(y)\in\mathbb{R}^{n}$ such that the RoPE attention logit approximates $f(d,x,y)$ with arbitrarily small error:
$\left|w(d,\bm{k}(x),\bm{q}(y))-f(d,x,y)\right|<\epsilon.$ (5)
3.3 Extending Discussions to General Relative Position Encoding
This decomposition naturally extends to ALiBi Press et al. (2021), another most common form of positional encoding in current LLMs (e.g., MPT model family Team (2023)). ALiBi introduces a position-dependent bias directly to the attention logits, which is linearly added and separable from the semantic interaction computed via the key-query dot product. This makes the position-semantic decomposition even more explicit and fits directly into the discussions above.
@@ 4 Position Generalization of LLMs
4200 5120 6144 7168 8192
KL (LM-Infinite) 0.0789 1.2093 2.6327 3.5192 4.4401
KL (Extending Vanilla LM) 0.5493 6.4862 8.6050 12.8479 15.9810
Log-Perplexity (LM-Infinite) 1.102 0.996 0.995 1.020 1.004
Log-Perplexity (Extending Vanilla LM) 1.316 7.268 8.737 8.754 8.901
Table 2: KL divergence between the distribution of extra-length features and pretraining-length features is significantly smaller when a length extrapolation technique (such as LM-Infinite) is applied compared to vanilla LLMs. The distribution was approximated with a multivariant Gaussian distribution. This trend aligns with the lower loss (lower log-perplexity) when applying a length extrapolation method.
@@ 4.1 Tolerance to Position Perturbations
The results, presented in Fig. 5, show that LLMs exhibit robustness to these perturbation methods. The original model has a perplexity of 3.688 on the ArXiv documents Gao et al. (2020a). Text transposition has a minimal impact on perplexity, with only a 0.02 increase when 1% of tokens are shuffled up to a distance of 1000 tokens. This suggests that LLMs do not rigidly depend on strict word order. Our intervention techniques simulate this linguistic transposition effect in Fig. 5(b) and  5(c). Feature transposition also introduces a modest increase in perplexity, indicating that while position indices contribute to contextual representation, their precise ordering is not always critical in each self-attention layer. As further analysis, when the position encoding contains perturbed indices, the perplexity still has a marginal increase when 10% of token positions were perturbed by ±$\pm 10$, or increase by an absolute value of 0.01 when 1% of tokens has position encoding perturbed by ±$\pm 10$. These phenomena further align with the previous observations that position information acts as a disentangled additive factor rather than being tightly entangled with semantic relationships.
@@ 4.2 When and How LLMs Generalize to Longer Texts
In vanilla relative position encoding, the list of distance values $[d]$ used in the attention operation between one query vector $q_{i}$ and all key features $k_{j}$ sequence is just the list of inter-token distances $[i,\ldots,0]$. Recent techniques like LM-Infinite and InfLLM enable LLMs to generalize to longer text sequences than those encountered during training. The common practice adopted by these techniques is to modify the relative position before applying the original self-attention mechanism. This is equivalent to applying self-attention over a modified $[{\bm{k}}_{i}]$ and $[{\bm{v}}_{i}]$ sequence, which might be different (sometimes significantly shorter) than the original $[{\bm{k}}_{i}]$ and $[{\bm{v}}_{i}]$ sequences. More specifically, in those techniques, the resulting sequence of features usually appears as if they are of the following positional distances:
@@
This is in contrast to the intuition we obtained from common machine learning practices: why do LLMs train purely from shorter texts that generalize to extreme lengths (e.g., 200M in LM-Infinite) with only minor modifications to the model architecture? Moreover, little explicit design was implemented in modern SotA LLMs to enable this extreme generalization. This phenomenon could find support in our analysis: even though posed to unseen extreme lengths, LLMs do not bind positional relevance information with the semantic features of the contextual tokens. In other words, the ${\bm{k}}$ and ${\bm{v}}$ vectors could be interpreted as a pool of semantic features. The self-attention mechanism approximately and additively applies the position component to the pool. As long as the resulting distance list is similar to its normal shape $[L_{PT},L_{PT}-1,\ldots,1,0]$, the attention output vector will reside in its normal distribution. Therefore, technically, the attention output vector is still in-distribution, so the remaining parts of LLMs on top of the attention outputs will not take out-of-distribution features as inputs.
To verify this claim, we visualize the attention output vectors of an arbitrary layer using the technique above on the Pile dataset Gao et al. (2020b). This is a projection down to a 2-D plane using PCA [^4] https://scikit-learn.org/stable/modules/generated/sklearn.decomposition.PCA.html. The blue dots are the normal attention output vectors, which mark their normal distribution. Then, we select a set of ${\bm{q}}$ vectors and associate them with different colors. For each vector, we apply it over a sub-sequence of length $L_{PT}$: $[{\bm{k}}_{i},{\bm{k}}_{i+1},\ldots,{\bm{k}}_{i+L_{PT}}]$ and $[{\bm{v}}_{i},{\bm{v}}_{i+1},\ldots,{\bm{v}}_{i+L_{PT}}]$. As we vary the value of starting position $i$, we trace the output vector with colored broken lines. As shown in Figure 6, these lines, though extending to different directions and different ranges depending on the ${\bm{q}}$, still wander within the range of normal attention output vector distribution.
In Table 2, we also empirically verify that, as the context length increases, the distribution of length-generalizable text features remains similar to normal feature distributions, as shown in the following table. However, vanilla models on longer lengths will have an increasing KL divergence from normal features. This trend aligns with the trend of final perplexity. This further validates our explanation of the length generalization and provides insights for future manipulation of the self-attention module for research purposes.
@@ 5 Conclusions and Future Work
In this work, we investigated the computational mechanisms behind the position generalization capabilities of LLMs. We first demonstrated that attention logits in LLMs can be approximately disentangled into independent components representing positional and semantic relevance. This finding suggests a structured decomposition within the model’s internal computations. Through empirical analysis, we further examined various forms of position generalization at the LLM representation level. These insights provide both computational explanations and insights into controlling these phenomena.
Future research could delve deeper into the specific architectural choices and training data patterns that contribute to this robustness. We also envision practical implications where the discovered decomposition might be important:
  • Improved efficiency in attention calculation: after estimating the semantic importance and positional patterns, some (regions of) attention logits can be estimated without calculating precisely. This will be especially useful when those logits are low so that estimation errors will only translate to a small deviation in their exponentials, which are used in the softmax operator later. Such techniques can be further incorporated into novel architectures, which will be inherently more efficient.
  • KV cache compression: Similar to the bullet above, the estimated importance component can be used as a precise and efficient criterion for filtering out $k,v$ caches that are more likely to receive low logits.
  • Position bias mitigation: LLMs are reported to favor options that appear at specific positions Wang et al. (2024) or are lost in the middle Liu et al. (2023). The discovered positional pattern and semantic importance can be used to re-calculate the attention logits, alleviating or removing such bias.
@@ Appendix D More Example Visualization from Other Models
Appendix E Proof of Theorem 1
Proof.
Let $\theta_{r}=2\pi\frac{r}{N}$, and let us view $f(d,x,y)$ along the variable of $d\in[0...L]$, and define the Fourier spectrum $F(x,y)=[F_{0}(x,y),...,F_{N-1}(x,y)]$:
$F_{r}(x,y):=\frac{1}{N}\sum_{d=0}^{L}f(d,x,y)e^{-\mathrm{i}d\theta_{r}}$
We now encode the spectrum into $\bm{k}(x)$ so that
$\mathrm{i}k_{2r}(x)+k_{2r+1}(x)=\frac{F_{r}(x,y)}{\mathrm{i}q_{2r}(y)+q_{2r+1}(y)}$
Then we have:
$w(d,\bm{k}(x),\bm{q}(y))$
$=$ $\sum_{r=1}^{N}\text{Re}\left[F_{r}(x,y)e^{id\theta_{r}}\right]$
$=$ $f(d,x,y)+O(e^{-cn})$
for some constant $c$ by established bounds of discrete (inverse) Fourier coefficients in Nissilä (2018). ∎